A sample of nitrogen occupies 5.50 liters under a pressure of 900 torr at 25oC. At what temperature will it occupy 10.0 liters at the same pressure? Show Complete Solution.


1
Expert's answer
2021-11-08T08:12:16-0500

V1/T1= V2/T2

5.5/298= 10/T2

0.0185T2= 10

T2= 10/0.0185

= 540.54K